OpenStudy (eric_d):

f(x)=x-2 find inverse function @ganeshie8

OpenStudy (neer2890):

to find inverse: let y=f(x) so y=x-2 now find x from above equation x=y+2 now interchange x and y so y=x+2 and inverse is f(x)=x+2

OpenStudy (eric_d):

okay

OpenStudy (eric_d):

f(x)= (x-2)^2 find inverse

OpenStudy (neer2890):

just follow the same procedure and let me know how you do it?

OpenStudy (eric_d):

let y = f(x) f^-1(y)=x y=(x-1)^2

OpenStudy (eric_d):

y^2=x-1 x=y^2+1

OpenStudy (eric_d):

is it correct

OpenStudy (neer2890):

where did you get x-1?

OpenStudy (eric_d):

frm question ... remove the power 2 then y becomes y^2

OpenStudy (eric_d):

@neer2890

OpenStudy (neer2890):

your question is f(x)=(x-2)^2 let y=f(x) y=(x-2)^2 squaring both sides y^2=x-2 x=y^2+2 now interchange x and y y=x^2+2 so invers is f(x)=x^2+2

OpenStudy (eric_d):

Thanks @neer2890